OpenStudy (anonymous):

Divide and simplify to lowest terms. 3x^2-12x/5x-3 / x-4/5x-3? i got.. Please let me know if this is correct. Thanks. 3x^2-12x/x-4

OpenStudy (anonymous):

i used casio classpad to simplify...idk if calculator is wrong or can't simplify further for strange reasons but i got...3x/(5x-3)^2

OpenStudy (anonymous):

i simplifies ur answer though....3x^2-12x/x-4 simplies further into 3x

OpenStudy (anonymous):

\[\frac{3x^2-12x}{5x-3}\times \frac{5x-3}{x-4} \] is the start

OpenStudy (anonymous):

the \(5x-3\) cancels leaving \[\frac{3x^2-12x}{x-4}\] \[=\frac{3x(x-4)}{x-4}=x\]

OpenStudy (anonymous):

so there's supposed to be multiply in between....k...then above answer is only half correct 3x^2-12x/x-4 becomes 3x

OpenStudy (anonymous):

so final is 3x

OpenStudy (anonymous):

so I didn't simplify it completely like I should have!

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thank you guys!!

OpenStudy (anonymous):

ur welcome